An Example for the Usage of the birkmult Class File

Birkhäuser Publishing Ltd. and A Second Author

To my boss

Abstract. The aim of this work is to provide the contributors to journals or to multi-authored books with an easy-to-use and flexible class file compatible with L^{A}TEX and AMS-L^{A}TEX.

Mathematics Subject Classification (2000). Primary 99Z99; Secondary 00A00.

Keywords. Class file, journal.

1. Document Preamble

Start the article with the command

\documentclass{birkmult}

After that, needed macro packages and new commands can be inserted as in every L^{A}TEX or AMS-L^{A}TEX document. Don’t use commands that change the page

layout (like \textwidth, \oddsidemargin etc.) or fonts.

Journals: for some journals, the citation lines can be included automatically by using the appropriate option in the above command, namely,

• Integral Equations and Operator Theory: \documentclass[ieot]{birkmult} • Journal of Mathematical Fluid Mechanics: \documentclass[jmfm]{birkmult} • Mediterranean Journal of Mathematics: \documentclass[medjm]{birkmult}

  • Milan Journal of Mathematics: \documentclass[mjm]{birkmult}

This work was completed with the support of our TEX-pert.

2. Frontmatter

The command

\begin{document}

starts – as always – the article.

2.1. Author Data Afterwards, insert title, author(s) and affiliation(s), as in the source file to this document, bmultdoc.tex. E.g.,

\title[An Example for birkmult]
 {An Example for the Usage of the\\ birkmult Class File}
%----------Author 1
\author[Birkh\"auser]{Birkh\"{a}user Publishing Ltd.}
\address{%
Viaduktstr. 42\\
P.O. Box 133\\
CH 4010 Basel\\
Switzerland}
\email{info@birkhauser.ch}

For each author the commands \author, \address and \email should be used separately. See the last page of this document for the typesetting layout of the above addresses.

2.2. Abstract, Thanks, Key Words, MSC The abstract environment typesets the abstract:

\begin{abstract}
The aim of this work is to provide the contributors to edited books
with an easy-to-use and flexible class file compatible with
\LaTeX\ and \AmS-\LaTeX.
\end{abstract}

In addition, the Mathematical Subject Codes, some key words and thanks can be given:

\thanks{This work was completed with the support of our
\TeX-pert.}
\subjclass{Primary 99Z99; Secondary 00A00}
\keywords{Class file, journal}

Finally, \maketitle typesets the title.

3. Mainmatter

Now type the article using the usual L^{A}TEX and (if you need them) AMS-L^{A}TEX commands.

We gratefully appreciate if the text does not contain \overfull and/or \underfull boxes, if equations do not exceed the indicated width, if hyphenations have been checked, and if the hierarchical structure of your article is clear. Please avoid caps and underlines.

Just to give examples of a few typical environments:

Definition 3.1. This serves as environment for definitions. Note that the text appears not in italics.

This is a sample equation: c^{2} = a^{2} + b^{2}  (3.1)

The above equation received the label testequation.

Theorem 3.2 (Main Theorem). In contrast to definitions, theorems appear typeset in italics as it has become more or less standard in most textbooks and monographs. Equations can be cited using the \eqref command which automatically adds brackets: \eqref{testequation} results in (3.1).

Proof. A special environment is predefined: the proof environment. Please use

\begin{proof}

proof of the statement

\end{proof}

for typesetting your proofs. The end-of-proof symbol □ will be added automatically. □

There are two known problems with the placement of the end-of-proof sign:

1. if your proof ends with a s i n g l e displayed line, the end-of-proof sign would be placed in the line below; if you want to avoid this, write your line in the form

    $$displayed math line \eqno\qedhere$$

which results in

Proof.

displayedmathline  □

2. if your proof ends with a larger displayed environment, you either may leave the end-of-proof sign where it is located or you could add some

    \vskip-1em

one line above the

    \end{proof}

command, in order to place the symbol a bit above.

Remark 3.3. Additional comments are being typeset without boldfaced entrance word as they may be minor important.

Example. For some constructs, even no number is required.

Displayed equations may be numbered like the following one:

\sqrt{1 − sin^{2}(x)} = | cos(x)|.  (3.2)

3.1. Here is a Sample Subsection Just needed because next thing is

3.1.1. Here is a Sample for a Subsubsection. One more sample will follow which clearly shows the difference between subsubsection deeper nested lists: Here is a Sample for a Paragraph. As you observe, paragraphs do not have numbers and start new lines after the heading, by default.

3.2. Indentation Though indentation to indicate a new paragraph is welcome, please do not use indentation when the new paragraph is already marked by an extra vertical space, as for example in the case of the first paragraph following a heading (this is standard in this class), or after using commands like \smallskip, \medskip, \bigskip etc.

3.3. Figures Please use whenever possible figures in EPS format (encapsulated postscript). Then, you can include the figure with the command

\includegraphics{figure.eps}

It is sometimes difficult to place insertions at an exact location in the final form of the article. Therefore, all figures and tables should be numbered and you should refer to these numbers within the text. Please avoid formulations like “the following figure… ”.

3.4. Your Own Macros If you prefer to use your own macros within your document(s) please don’t forget to send them to us together with the source files for the manuscript. We will need all these files to produce the final layout.

4. Backmatter

At the end of the document, the affiliation(s) will be typeset automatically. For this it is necessary that you used the \address command for including your affiliation, as explained above.

References

[1] A. B. C. Test, On a Test. J. of Testing 88 (2000), 100–120.

[2] G. Grätzer, Math into L^{A}TEX. 3rd Edition, Birkhäuser, 2000.

An Example for birkmult 5

Acknowledgment Many thanks to our TEX-pert for developing this class file.

Birkhäuser Publishing Ltd. Viaduktstr. 42 P.O. Box 133 CH 4010 Basel Switzerland

e-mail: info@birkhauser.ch

A Second Author The address of the second author sitting somewhere in the world

e-mail: dont@know.who.knows